OpenStudy (anonymous):

can someone please remind this nerd mom how to solve this 6-x-3=10

OpenStudy (anonymous):

first, combine like terms on the left side of the equation

OpenStudy (anonymous):

\[6-x-3=10\] \[3-x=10\]

OpenStudy (anonymous):

Then subtract 3 from both sides

OpenStudy (anonymous):

you end up with -x=7

OpenStudy (anonymous):

so what would it look like written out in the solution steps? Its been a while since I was in school...and my brain needs refreshed :)

OpenStudy (anonymous):

6-x-3 = 10 switch position of -x and -3 6-3-x=10 combine like terms (6-3)=3 so 3 - x= 10 (subtract 9 from both sides to move 3 from left to right) -3 -3 0 - x= 7 now we need to divide both sides by (-1) \[\frac{ -x }{ -1 }=\frac{7}{-1}\] x=-7

OpenStudy (anonymous):

Thank you!!